# Nonlocal interaction and quantum friction in sliding Bi2Se3 topological surfaces

### v3

@@ Abstract
Topological insulator Bi2Se3 thin films exhibit unique electronic properties arising from their topologically protected surface states. In a theoretical model capturing the essential physics of Dirac electrons in Bi2Se3, we study excitations and dissipation in two infinite parallel metallic plates undergoing relative motion. The degrees of freedom of the electrons in both plates are modeled using the 1+2 dimensional Dirac field, and a nonlocal potential is selected to describe the interaction between the two plates. The internal relative motion is introduced via a Galilean boost, with one plate assumed to slide relative to the other. We then calculate the effective action of the system and derive the vacuum occupation number in momentum space using a perturbative method. Numerical plots reveal that the vacuum occupation number, as a function of momentum, is isotropic for a motion speed $v=0$ and anisotropic for nonzero $v$. The relative motion induces energy transfer between the plates, leading to on-shell excitations in a manner analogous to the dissipative process of the Schwinger effect. Consequently, we study the motion-induced dissipation effects and the dissipative forces through the quantum action. By using experimental Fermi velocities of Bi2Se3, our results demonstrate that both the imaginary part of the quantum action due to the motion boost and the dissipative force exhibit a threshold as functions of $v$, and both are positively correlated with $v$.
@@ 1 Introduction
Topological insulators represent a novel quantum state of matter characterized by an insulating bulk and topologically protected gapless surface states [1, 2]. Among various topological insulator materials, Bi2Se3 has emerged as an ideal platform due to its relatively large bulk band gap and simple surface electronic structure [3, 4]. Particularly in the thin-film limit, Bi2Se3 exhibits remarkable tunability of its topological surface states: surface state hybridization opens a tunable gap at the Dirac point [5], surface engineering enables optimization of Dirac cone isolation [6], and strain can modulate the bulk gap and even induce topological phase transitions [7]. These highly controllable topological interface states provide a unique platform for investigating fundamental physical processes in low-dimensional quantum systems, especially in exploring quantum dissipation mechanisms induced by relative motion, which is of paramount importance for understanding energy transfer in nanoscale devices.
@@
In this paper, we study the quantum excitations and dissipation of two infinite parallel metallic plates moving relative to each other along their interface. This system is similar to the 3D topological insulator $Bi_{2}Se_{3}$ [32, 3]. Massive Dirac electrons are considered in this model and a nonlocal Anderson mixed coupling is introduced to transmit friction, instead of relying on vacuum fluctuations of the electromagnetic field. We assume a relative sliding is driven by an external field for excitations and dissipation, and analyze the interaction of electrons in the two plates via a nonlocal potential. To study the excitations, we start with the DOFs associated with each plate using massive Dirac fermions and write the effective action of the electrons by considering one plate sliding relative to the other. By focusing on the time evolution of the system, we calculate the probability of particle production and derive the vacuum occupation number of the fixed plate perturbatively. Since the process of relative motion inducing on-shell excitations is very similar to the Schwinger effect, we can study the quantum effects of this process through the method of path integrals, and then obtain the imaginary part of the quantum action and the dissipative force, with real experimental Fermi velocity $v_{\mathrm{F}}$.
@@
The main structure of this paper is as follows. In Section 2, we introduce the electronic structures and topological properties of Bi2Se3 that motivate our theoretical framework. In Section 3, the microscopic model will be given. Section 4 discuss the time evolution of the system and calculatethe probability of particle production. In Section 5, we focus on the excitations of the system. Section 6 is about quantum effects and dissipation. Section 7 is the conclusion of this work.
2 Electronic structures and topological properties of three-dimensional and two-dimensional $Bi_{2}Se_{3}$
Figure 1: The band structure of Bi2Se3 in (a) three-dimension and (b) two-dimension.
The theoretical model developed in this study investigates motion-induced dissipation between two parallel plates described by massive Dirac fermions. The model is motivated by the distinctive electronic properties of Bi2Se3 [32, 3], which band structure is given in Figure 1. The low-energy excitations on the surface of Bi2Se3 are governed by massless Dirac fermions that exhibit spin–momentum locking, a fundamental feature incorporated into the Hamiltonian of our model. Consequently, understanding the electronic structure of Bi2Se3 in both its three-dimensional bulk form and its two-dimensional thin-film limit is essential for establishing the theoretical framework of this work.
Bismuth selenide (Bi2Se3) crystallizes in a rhombohedral structure with space group $R\bar{3}m$ (No. 166). The crystal can be conveniently described as a stack of quintuple layers (QLs) along the trigonal $c$ axis, where each QL consists of atomic planes arranged in the sequence Se(1)–Bi–Se(2)–Bi–Se(1). The bonding within each QL is predominantly covalent–ionic, whereas adjacent QLs are held together by weak van der Waals interactions. This layered configuration forms the structural basis for thin-film exfoliation and epitaxial growth [5, 6]. As a direct consequence of the bulk topology, the material hosts robust, gapless metallic surface states that can be described by a massless Dirac Hamiltonian. These surface states form a Dirac cone with a linear energy dispersion $E(k)=\pm\hbar v_{\mathrm{F}}|\bm{k}|$, where $v_{F}$ denotes the Fermi velocity. The electron spin is oriented perpendicular to its momentum, a phenomenon known as spin–momentum locking, which suppresses backscattering from nonmagnetic impurities and preserves the topological protection of the surface states [5, 7].
When the thickness of Bi2Se3 is reduced to the nanometer scale, typically below about six quintuple layers, a tunable dimension arises due to quantum confinement and intersurface coupling. In this quasi-two-dimensional configuration, the topological surface states from the top and bottom surfaces, which are independent in thick crystals, begin to hybridize. This hybridization opens an energy gap $\Delta_{f}$ at the Dirac point [5]. The magnitude of this gap is highly sensitive to the film thickness $d$, offering an effective means to tailor the electronic structure. This tunable energy gap plays a central role in the present model, as it directly influences the density of low-energy states and the phase space available for excitations. The Dirac cone of Bi2Se3 can be effectively tuned through surface engineering, such as oxygen substitution on surface Se sites, which modifies the surface potential and isolates the Dirac cone within the bulk band gap, thereby enhancing topological transport [6]. In addition, the electronic structure is highly sensitive to external strain, where uniaxial compression can alter the bulk band gap and even induce a topological phase transition by modifying spin–orbit coupling and Coulombic interactions [7].
In this study, the essential physics of two coupled Bi2Se3 surfaces is modeled by two parallel slabs whose internal degrees of freedom are described by a (2+1)-dimensional massive Dirac field. The massive Dirac Hamiltonian used in our analysis (Eq. (3)) directly corresponds to the effective Hamiltonian of Bi2Se3 surface states. The nonlocal interaction potential $U(x,y)$ between the slabs (Eq. (5)) represents the intersurface hybridization in Bi2Se3 films, or more generally, the coupling between two low-dimensional Dirac systems. The relative motion of the slabs introduces a Galilean acceleration, creating a nonequilibrium state analogous to that between two topological insulator surfaces. The relative velocity $v$ serves as a tuning parameter, similar to strain, thickness, or chemical substitution that modify the properties of real two-dimensional Bi2Se3 structures. The anisotropy of the vacuum occupation number $n_{L}(\bm{k})$ caused by a nonzero $v$ (Section 5) and the resulting quantum dissipation (Section 6) can be experimentally probed through sliding interfaces or nonequilibrium transport in topological insulator films and heterostructures. The tunable electronic structure of both three-dimensional and two-dimensional Bi2Se3 therefore provides a robust physical foundation and a promising materials platform for studying the quantum dissipation phenomena discussed in this work.
@@ 3 The microscopic model
The model of Bi2Se3 system we consider in this study is the interface between two infinite metallic plates, which is shown in Fig. 2. For easy identification, one plate is marked as L-plate and the other as R-plate. Since our purpose is to study motion driven excitations and dissipation, relative sliding is required in the system. Here we assume that the L-plate is at rest in the laboratory frame, and the R-plate is moving along the interface. The total area of each plate is $V$, and the total evolution time for fiction of the system is $T$. To analyze the model, a 1+3 dimensional space-time coordinate system is constructed as $X=\begin{pmatrix}x^{0},x^{1},x^{2},x^{3}\end{pmatrix}^{\mathsf{T}}$, where $x^{0}=ct$ is the time coordinate with the speed of light $c$ and $\{x^{1},x^{2},x^{3}\}$ are the Cartesian coordinates. The axis of $x^{3}$ is set to be perpendicular to the L-plate. The space-time coordinates for the 1+2 dimenstional space-time domain of the L-plate read $x=\begin{pmatrix}x^{0},x^{1},x^{2}\end{pmatrix}^{\mathsf{T}}$ and the spacial components are written as $\emph{{x}}=\begin{pmatrix}x^{1},x^{2}\end{pmatrix}^{\mathsf{T}}$. We follow the convention of natural unit such that $\hbar=c=1$. The corresponding covariant components of a space time vector is $k_{\mu}=\eta_{\mu\nu}k^{\nu}$, i.e. $k_{0}=k^{0}$, $k_{1}=-k^{1}$ and $k_{2}=-k^{2}$, where $\eta_{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$ is the Minkovski metric. So the inner product of two space time vectors is $k\cdot x=k^{0}x^{0}-\emph{{k}}\cdot\emph{{x}}$.
@@
In this study of Bi2Se3 model, electrons on the two plates are considered to couple via a direct-exchange interaction, similar to the Anderson mixed model [34]. It involves local interactions and the electron-electron interaction typically involves direct exchange, ignoring nonlocal electron exchange coupling. However, the interaction occurs between two many-electron systems with nonlocal coupling, it should include not only a direct exchange term but also a nonlocal term. By taking a real-valued nonlocal potential $U(x,y)=U(x^{0},\emph{{x}};y^{0},\emph{{y}})$, this nonlocal Anderson mixed model where the fermions belonging to the two plates interact with each other can be written as:
@@ 5 Excitations of L-plate
As a function of momentum, $n_{\mathrm{L}}(\emph{{k}})$ is isotropic for $v=0$, but becomes anisotropic for nonzero $v$. Moreover, $n_{\mathrm{L}}(\emph{{k}})$ is time-independent, indicating that our approximation applies to a NESS regime where the system remains close to the initial vacuum state. The nonzero vacuum occupation number in momentum space shows that the interaction modifies the free vacuum, and the relative motion induces on-shell excitations from the vacuum.
@@ 6 Quantum dissipation
Here we derive the expression for the imaginary part of the quantum action and the dissipative force between the two plates of Bi2Se3 surfaces. In our model, the R-plate is subject to an external force, rendering it an open system. Consequently, the vacuum state of the system becomes unstable, leading to particle excitations within the plates. Therefore, dissipation is associated with the imaginary part of the quantum action $\Gamma$. The quantum transition amplitude has a physical interpretation: its square represents the probability of the corresponding transition process. In this study, the probability of the vacuum-to-vacuum transition process is expressed as $|Z|^{2}$. For small $\mathrm{Im}\Gamma$, this probability can be approximated perturbatively as:
@@
$\mathrm{Im}\Gamma=\frac{g^{2}v_{\mathrm{F}}VT}{4}(v^{2}-4v_{\mathrm{F}}^{2})^{\frac{3}{2}}\int\frac{\mathrm{d}k^{2}}{2\pi}\frac{1}{\sqrt{v_{\mathrm{F}}^{2}(k^{2})^{2}+v_{\mathrm{F}}^{2}m^{2}}}\frac{1}{(1-9v_{\mathrm{F}}^{2})v^{2}(k^{2})^{2}-(9v^{2}-4)v_{\mathrm{F}}^{2}m^{2}}.$ (51)
Now we substitute the experimentally measured values of the Fermi velocity $v_{\mathrm{F}}$ into Eq. (51) to examine how $\mathrm{Im}\Gamma$ varies with $v_{\mathrm{F}}$. Three representative Fermi velocities reported for Bi2Se3 are considered. The earliest angle-resolved photoemission spectroscopy (ARPES) measurements of the surface-state dispersion yield $v_{\mathrm{F}}\approx 3\times 10^{5}$ m/s [39, 40]. A slightly larger value, $v_{\mathrm{F}}\approx 4.2\times 10^{5}$ m/s, was obtained from ARPES measurements of the linear dispersion of the topological surface states $E(k)$, where the slope was determined using $E=\hbar v_{\mathrm{F}}k$ [41, 42, 43]. The third value, $v_{\mathrm{F}}\approx 2.2\times 10^{5}$ m/s, was extracted from Shubnikov–de Haas (SdH) quantum oscillation data for pure Bi2Se3 [44]. Figure 3 presents the numerical results of Eq. (51) for these three values of $v_{\mathrm{F}}$ with $m=1$ fixed, showing a threshold at the relative velocity $v=2v_{\mathrm{F}}$. This result indicates that dissipation occurs when $v>2v_{\mathrm{F}}$. The parameter $m$, which represents a controllable external magnetic field, has little influence on the overall trend of $\mathrm{Im}\Gamma$ as a function of $v_{\mathrm{F}}$ and only causes a rescaling of the coordinate axes. We give the test of different values of $m$ for Eq. (51) in Appendix A.
Figure 3: The imaginary part of quantum action as a function of $v$, for the case $v_{\mathrm{F}}=3\times 10^{5}$$4.2\times 10^{5}$ and $2.2\times 10^{5}$, with $m=1$, in units of $g^{2}TV$.
@@
$F_{\mathrm{diss}}=\frac{3g^{2}v_{\mathrm{F}}}{4v}(v^{2}-4v_{\mathrm{F}}^{2})^{\frac{3}{2}}\int\frac{\mathrm{d}k^{2}}{2\pi}\frac{1}{\sqrt{v_{\mathrm{F}}^{2}(k^{2})^{2}+v_{\mathrm{F}}^{2}m^{2}}}\frac{1}{(1-9v_{\mathrm{F}}^{2})v^{2}(k^{2})^{2}-(9v^{2}-4)v_{\mathrm{F}}^{2}m^{2}}.$ (56)
Similarily, we use the three values of $V_{\mathrm{F}}$ for $\mathrm{Im}\Gamma$ to demostrate the numerical result of $F_{\mathrm{diss}}$ with $m=1$ in Fig. 4. We can see that it is a threshold of the relative velocity $v=2v_{\mathrm{F}}$, like the $v$-dependence of $\mathrm{Im}\Gamma$ in Fig. 3. The difference is, the in Fig. 4 the curve of $F_{\mathrm{diss}}$ shows linear behaviors, and the slope of the $v-F_{\mathrm{diss}}$ curves give linear frictions coefficient at zero temperature.
Figure 4: The dissipative force as a function of $v$, for the case $v_{\mathrm{F}}=3\times 10^{5}$$4.2\times 10^{5}$ and $2.2\times 10^{5}$, with $m=1$, in units of $g^{2}TV$.
@@
Our results indicate that under relative motion, both the imaginary part of the quantum action $\mathrm{Im}\Gamma$ and the dissipative force $F_{\mathrm{diss}}$ exhibit a distinct velocity threshold behavior. When the relative velocity $v$ exceeds twice the Fermi velocity $2v_{\mathrm{F}}$, the system undergoes pronounced quantum dissipation. This threshold provides a clear criterion for experimentally detecting and controlling energy dissipation between Bi2Se3 interfaces. Moreover, a linear relationship between the dissipative force and the relative velocity is observed, implying the existence of a linear friction coefficient at zero temperature. This finding offers new insights into the energy-loss mechanisms of Bi2Se3 films under sliding contact or nonequilibrium transport. Overall, these results deepen the understanding of quantum dissipation in topological insulators and provide a theoretical foundation for designing low-dissipation nanodevices and developing quantum interface control strategies based on Bi2Se3.
@@ 7 Conclusion
In this paper, we studied the excitation and quantum dissipation of two Bi2Se3 surfaces induced by the internal relative motion of two parallel metallic plates. The degrees of freedom (DOFs) of the electrons in both plates were modeled using the 1+2 dimensional Dirac field, and a nonlocal potential was chosen to describe the interaction between L-electrons and R-electrons. The internal relative motion was introduced via a Galilean boost.
@@
Subsequently, we derived the expression for the imaginary part of the quantum action $\mathrm{Im}\Gamma$ due to the relative motion and calculated the dissipative force $F_{\mathrm{diss}}$ between the two plates. The numerical results show that both $\mathrm{Im}\Gamma$ and $F_{\mathrm{diss}}$ exhibit a threshold and are positively correlated with $v$. Notably, the dependence of $F_{\mathrm{diss}}$ on $v$ follows a linear relationship, leading to a linear friction coefficient at zero temperature. The dissipation velocity threshold ($v>2v_{\mathrm{F}}$) and the linear friction relationship revealed in this study provide a solid theoretical foundation for experimentally detecting and controlling quantum energy dissipation at Bi2Se3 interfaces. These findings are crucial for understanding the nonequilibrium transport behavior of this topological insulator in nanoscale systems and for guiding the design of low-dissipation quantum devices.
In this study, the dissipation rate threshold ($v>2v_{\mathrm{F}}$) and the linear friction relationship revealed by the theoretical model of Bi2Se3 provide a solid theoretical foundation for experimentally detecting and controlling quantum energy dissipation between its topological interfaces. It is worth highlighting that we employed a specific physical motivation for the nonlocal interaction $U$ in our calculations. The coupling between the two plates is assumed to be s-d coupling, which originates from fundamental electromagnetic interactions and is inherently nonlocal. Since condensed matter systems often feature complex and novel excitations, this method could also be applied to other systems beyond electrons. However, for strongly correlated systems without a well-defined quasi-particle picture, non-perturbative approaches to quantum dissipation are also worth exploring.
Appendix A The impact of different values $m$ on $\mathrm{Im}\Gamma$ and $F_{\mathrm{diss}}$
Here we discuss the changes in the results of $\mathrm{Im}\Gamma$ and $F_{\mathrm{diss}}$ under different values of $m$. While keeping the value of $v_{F}$ unchanged and substituting $m=0.01$ into $\mathrm{Im}\Gamma$ (Eq. (51)), Fig. 8(a) shows that its magnitude decreases along the y-axis as $m$ increases, while the shapes of the three curves remain unchanged. The similar behavior is observed for $F_{\mathrm{diss}}$ (Eq. (56)) in Fig. 8(b), indicating that in our model, variations in the external magnetic moment affect only the magnitudes of $\mathrm{Im}\Gamma$ and $F_{\mathrm{diss}}$.

### v5

# Motion-driven quantum dissipation in an open electronic system with nonlocal interaction

@@ Abstract
In this paper, we study excitations and dissipation in two infinite parallel metallic plates undergoing relative motion. The degrees of freedom of the electrons in both plates are modeled using the 1+2 dimensional Dirac field, and a nonlocal potential is selected to describe the interaction between the two plates. The internal relative motion is introduced via a Galilean boost, with one plate assumed to slide relative to the other. We then calculate the effective action of the system and derive the vacuum occupation number in momentum space using a perturbative method. Numerical plots reveal that the vacuum occupation number, as a function of momentum, is isotropic for a motion speed $v=0$ and anisotropic for nonzero $v$. The relative motion induces energy transfer between the plates, leading to on-shell excitations in a manner analogous to the dissipative process of the Schwinger effect. Consequently, we study the motion-induced dissipation effects and the dissipative forces through the quantum action. Numerical results demonstrate that both the imaginary part of the quantum action due to the motion boost and the dissipative force exhibit a threshold as functions of $v$, and both are positively correlated with $v$.
@@ 1 Introduction
In this paper, we study the quantum excitations and dissipation of two infinite parallel metallic plates moving relative to each other along their interface. This system is similar to the 3D topological insulators, such as $Bi_{2}Se_{3}$, $Sb_{2}Te_{3}$ and $Bi_{2}Te_{3}$ [25, 26]. Massive Dirac electrons are considered in this model and a nonlocal Anderson mixed coupling is introduced to transmit friction, instead of relying on vacuum fluctuations of the electromagnetic field. We assume a relative sliding is driven by an external field for excitations and dissipation, and analyze the interaction of electrons in the two plates via a nonlocal potential. To study the excitations, we start with the DOFs associated with each plate using massive Dirac fermions and write the effective action of the electrons by considering one plate sliding relative to the other. By focusing on the time evolution of the system, we calculate the probability of particle production and derive the vacuum occupation number of the fixed plate perturbatively. Since the process of relative motion inducing on-shell excitations is very similar to the Schwinger effect, we can study the quantum effects of this process through the method of path integrals, and then obtain the imaginary part of the quantum action and the dissipative force.
@@
The main structure of this paper is as follows. In Section 2, the microscopic model will be given. Section 3 discuss the time evolution of the system and calculatethe probability of particle production. In Section 4, we focus on the excitations of the system. Section 5 is about quantum effects and dissipation. Section 6 is the conclusion of this work.
@@ 2 The microscopic model
The model of system we consider in this study is the interface between two infinite metallic plates, which is shown in Fig. 1. For easy identification, one plate is marked as L-plate and the other as R-plate. Since our purpose is to study motion driven excitations and dissipation, relative sliding is required in the system. Here we assume that the L-plate is at rest in the laboratory frame, and the R-plate is moving along the interface. The total area of each plate is $V$, and the total evolution time for fiction of the system is $T$. To analyze the model, a 1+3 dimensional space-time coordinate system is constructed as $X=\begin{pmatrix}x^{0},x^{1},x^{2},x^{3}\end{pmatrix}^{\mathsf{T}}$, where $x^{0}=ct$ is the time coordinate with the speed of light $c$ and $\{x^{1},x^{2},x^{3}\}$ are the Cartesian coordinates. The axis of $x^{3}$ is set to be perpendicular to the L-plate. The space-time coordinates for the 1+2 dimenstional space-time domain of the L-plate read $x=\begin{pmatrix}x^{0},x^{1},x^{2}\end{pmatrix}^{\mathsf{T}}$ and the spacial components are written as $\emph{{x}}=\begin{pmatrix}x^{1},x^{2}\end{pmatrix}^{\mathsf{T}}$. We follow the convention of natural unit such that $\hbar=c=1$. The corresponding covariant components of a space time vector is $k_{\mu}=\eta_{\mu\nu}k^{\nu}$, i.e. $k_{0}=k^{0}$, $k_{1}=-k^{1}$ and $k_{2}=-k^{2}$, where $\eta_{\mu\nu}=\mathrm{diag}(1,-1,-1,-1)$ is the Minkovski metric. So the inner product of two space time vectors is $k\cdot x=k^{0}x^{0}-\emph{{k}}\cdot\emph{{x}}$.
@@
In this study, electrons on the two plates are considered to couple via a direct-exchange interaction, similar to the Anderson mixed model [28]. It involves local interactions and the electron-electron interaction typically involves direct exchange, ignoring nonlocal electron exchange coupling. However, the interaction occurs between two many-electron systems with nonlocal coupling, it should include not only a direct exchange term but also a nonlocal term. By taking a real-valued nonlocal potential $U(x,y)=U(x^{0},\emph{{x}};y^{0},\emph{{y}})$, this nonlocal Anderson mixed model where the fermions belonging to the two plates interact with each other can be written as:
@@ 4 Excitations of L-plate
The numerical results of the vacuum occupation number for $v_{\mathrm{F}}=0.001$ and $m=10$ in the cases of $v=0$ and $v=0.7$ re shown in Fig. 2 as examples. Comparing subfigures in the top row ($v=0$) with those in the bottom row ($v=7$), it is evident that as $v$ increases, the occupation number is stretched along the $k_{1}$-direction, corresponding to the variation of $v$ along the $x^{1}$-axis. This result is expected because the relative motion is along the $x^{1}$-axis, meaning an external force acts on the R-plate in that direction, which causes the momentum of the L-electrons to vary accordingly. As a function of momentum, $n_{\mathrm{L}}(\emph{{k}})$ is isotropic for $v=0$, but becomes anisotropic for nonzero $v$. Moreover, $n_{\mathrm{L}}(\emph{{k}})$ is time-independent, indicating that our approximation applies to a NESS regime where the system remains close to the initial vacuum state. The nonzero vacuum occupation number in momentum space shows that the interaction modifies the free vacuum, and the relative motion induces on-shell excitations from the vacuum.
(a) 3D-plot, $v=0$
(b) Contour-plot, $v=0$
(c) 3D-plot, $v=7$
(d) Contour-plot, $v=7$
Figure 2: 3D and Contour-plot of the vacuum occupation number, for the case $v_{\mathrm{F}}=0.001$, $m=10$, in the order of $g^{2}$. Here (a)  and (b) for $v=0$; (c) and (d) for $v=7$.
@@ 5.1 The quantum action
5.2 The imaginary part of the quantum action and the dissipative force
In this section, we derive the expression for the imaginary part of the quantum action and the dissipative force between the two plates. In our model, the R-plate is subject to an external force, rendering it an open system. Consequently, the vacuum state of the system becomes unstable, leading to particle excitations within the plates. Therefore, dissipation is associated with the imaginary part of the quantum action, $\Gamma$. The quantum transition amplitude has a physical interpretation: its square represents the probability of the corresponding transition process. In this study, the probability of the vacuum-to-vacuum transition process is expressed as $|Z|^{2}$. For small $\mathrm{Im}\Gamma$, this probability can be approximated perturbatively as:
@@ 5.2 The imaginary part of the quantum action and the dissipative force
$\mathrm{Im}\Gamma=\frac{g^{2}v_{\mathrm{F}}VT}{4}\sqrt{v^{2}-4v_{\mathrm{F}}^{2}}\int\frac{\mathrm{d}k^{2}}{2\pi}\frac{1}{\sqrt{v_{\mathrm{F}}^{2}(k^{2})^{2}+v_{\mathrm{F}}^{2}m^{2}}}\frac{1}{(1-9v_{\mathrm{F}}^{2})v^{2}(k^{2})^{2}-(9v^{2}-4)v_{\mathrm{F}}^{2}m^{2}}.$ (51)
In Fig. 3, we give the numerical result of Eq. (51) for $v_{\mathrm{F}}=0.001$ and $m=10$, which shows a threshold of the relative velocity $v=2v_{\mathrm{F}}=0.002$. This indicates that dissipation occurs for $v>2v_{\mathrm{F}}$.
Figure 3: The imaginary part of quantum action as a function of $v$, for the case $v_{\mathrm{F}}=0.001$, $m=10$, in units of $g^{2}TV$.
@@
$F_{\mathrm{diss}}=\frac{3g^{2}v_{\mathrm{F}}}{4v}\sqrt{v^{2}-4v_{\mathrm{F}}^{2}}\int\frac{\mathrm{d}k^{2}}{2\pi}\frac{1}{\sqrt{v_{\mathrm{F}}^{2}(k^{2})^{2}+v_{\mathrm{F}}^{2}m^{2}}}\frac{1}{(1-9v_{\mathrm{F}}^{2})v^{2}(k^{2})^{2}-(9v^{2}-4)v_{\mathrm{F}}^{2}m^{2}}.$ (56)
Fig. 4 shows the numerical result of $F_{\mathrm{diss}}$ for $v_{\mathrm{F}}=0.001$ and $m=10$. We can see that it is a threshold of the relative velocity $v=2v_{\mathrm{F}}=0.002$, like the $v$-dependence of $\mathrm{Im}\Gamma$ in Fig. 3. The difference is, the in Fig. 4 the curve of $F_{\mathrm{diss}}$ shows a linear behavior, and the slope of the $v-F_{\mathrm{diss}}$ curve gives a linear friction coefficient at zero temperature.
Figure 4: The dissipative force as a function of $v$, for the case $v_{\mathrm{F}}=0.001$, $m=10$, in units of $g^{2}$.
@@ 6 Conclusion
In this paper, we studied the excitation and quantum dissipation induced by the internal relative motion of two parallel metallic plates. The degrees of freedom (DOFs) of the electrons in both plates were modeled using the 1+2 dimensional Dirac field, and a nonlocal potential was chosen to describe the interaction between L-electrons and R-electrons. The internal relative motion was introduced via a Galilean boost.
@@
Subsequently, we derived the expression for the imaginary part of the quantum action $\mathrm{Im}\Gamma$ due to the relative motion and calculated the dissipative force $F_{\mathrm{diss}}$ between the two plates. The numerical results show that both $\mathrm{Im}\Gamma$ and $F_{\mathrm{diss}}$ exhibit a threshold and are positively correlated with $v$. Notably, the dependence of $F_{\mathrm{diss}}$ on $v$ follows a linear relationship, leading to a linear friction coefficient at zero temperature.
It is worth highlighting that we employed a specific physical motivation for the nonlocal interaction $U$ in our calculations. The coupling between the two plates is assumed to be s-d coupling, which originates from fundamental electromagnetic interactions and is inherently nonlocal. Since condensed matter systems often feature complex and novel excitations, this method could also be applied to other systems beyond electrons. However, for strongly correlated systems without a well-defined quasi-particle picture, non-perturbative approaches to quantum dissipation are also worth exploring.